OpenStudy (anonymous):

HELP! Find the solution of the differential equation (1+x^8)y'+8x^7y=2e^2x , such that y(0)=0

OpenStudy (anonymous):

Solve the corresponding homogenous equation first. \[\Large (1+x^8)y'+8x^7y=0 \] Rearrange it to get: \[\Large y'+ \frac{8x^7}{1+x^8}y=0 \]

OpenStudy (anonymous):

So your integration factor is pretty obvious here, \[\Large \mu(x)=1+x^8 \] In fact, you can apply this to the general form of the differential equation, it's not necessary to solve the homogenous first.

OpenStudy (anonymous):

Rather a matter of art I would say, depending if one rather prefers to use Variation of Constants or partial integration.

OpenStudy (jamesj):

Notice that (d/dx)(1 + x^8) = 8x^7 hence you write the left hand side as \[ \frac{d\ }{dx}[(1+x^8)y] = (1+x^8)y' + 8x^7y\] and therefore the equation can be written as \[ \frac{d\ }{dx}[(1+x^8)y] = 2e^{2x} \] Now integrate both sides

OpenStudy (anonymous):

ok thanks!